# Universal scaling limits for spin networks via martingale methods

### v1

@@ 3.3 Example 1: Curie–Weiss model
$H_{ij}=\beta J_{0}/N,\ \ \ g\left(V\right)=N$ (97)
Since the prescription for the existence of the thermodynamic limit is that the components of the cavity fields stay finite as the system grows, then the cavity field must be normalised by a quantity that is linear in the number of spin. Then we must have the following form of the cavity field
@@
Most important, we can explicitly compute the coefficients
$\hat{\gamma}_{\ell}=p_{\ell},\ \ \ \bar{\gamma}_{\ell}=q_{\ell-1},\ \ \ \hat{\gamma}_{\ell}/\bar{\gamma}_{\ell}=p_{\ell}/q_{\ell}\rightarrow 0$ (100)
@@
$\bar{H}_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)=NJ_{0}\,q_{\ell-1}m\left(\sigma_{Q_{\ell-1}}\right)=NJ_{0}\,\sum_{\ell^{*}<\ell}p_{\ell^{*}}m\,(\sigma_{V_{\ell^{*}}})$ (102)
Notice in the last step we expanded the magnetisation states into the local magnetisations: we used a different superscript for the auxiliary variable $\ell^{*}$ to avoid confusion with the sub-index $\ell^{{}^{\prime}}$ before that has quite different meaning. The free energy functional of the blanket is then
$\bar{f}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)=-\frac{1}{\beta}\log 2\cosh\,[\,\beta J_{0}\,q_{\ell-1}m\,(\sigma_{Q_{\ell-1}})]=\\ =-\frac{1}{\beta}\log 2\cosh\,\beta J_{0}\sum_{\ell^{*}<\ell}p_{\ell^{\prime}}m\,(\sigma_{V_{\ell^{*}}})$ (103)
@@ 3.4 Example 2: Sherrington–Kirckpatrick model
the normalisation is the square root of the number of spins
this immediately imply the following from of the coefficients
$\hat{\gamma}_{\ell}=\sqrt{p_{\ell}},\ \ \ \bar{\gamma}_{\ell}=\sqrt{q_{\ell-1}},\ \ \ \hat{\gamma}_{\ell}/\bar{\gamma}_{\ell}=\sqrt{p_{\ell}/q_{\ell}}\rightarrow 0$ (107)
@@
$\hat{H}_{\ell}\left(\sigma_{V_{\ell}}\right)=\sqrt{p_{\ell}}\ H_{\mathrm{SK}}\left(\sigma_{V_{\ell}}\right),\ \ \ \bar{H}_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)=\sigma_{V_{\ell}}\cdot\sqrt{q_{\ell-1}}\ h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)$ (108)
@@
$\sqrt{q_{\ell-1}}\ h_{i}(\sigma_{Q_{\ell-1}})=\sum_{\ell^{*}<\ell}\sqrt{p_{\ell^{*}}}\,\,h_{i}(\sigma_{V_{\ell^{*}}})$ (109)
@@
$\bar{f}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)=-\frac{1}{\beta|V_{\ell}|}\sum_{i\in V_{\ell}}\log 2\cosh\,\beta\sqrt{q_{\ell-1}}\,\,h_{i}(\sigma_{Q_{\ell-1}})=\\ =-\frac{1}{\beta|V_{\ell}|}\sum_{i\in V_{\ell}}\log 2\cosh\,\beta\sum_{\ell^{*}<\ell}\sqrt{p_{\ell^{*}}}\,\,h_{i}(\sigma_{V_{\ell^{*}}})$ (110)
As said before, an alternative way to find the functional is the Cavity Method, [4, 33]: we showed a detailed derivation for the SK model in [1, 2, 3] (see also ref. [34] for the shortest derivation). Also, look at the Section 4, paragraph 4.2.1 of [10] for how to obtain the RSB theory as special (stationary) case of the kernel picture, within the larger context of Lattice Field Theories [3, 10, 11, 19, 20].
@@ Acknowledgements
We are grateful to Pan Liming (USTC), Giampiero Bardella (Sapienza Università di Roma) and Nicola Kistler (Goethe University Frankfurt) for interesting discussions and suggestions. This research was partially funded by the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 694925).

### v2

@@ 3.3 Example 1: Curie–Weiss model
$H_{ij}=\beta J_{0}/N,\ \ \ g\left(V^{2}\right)=N$ (97)
Since the prescription for the existence of the thermodynamic limit is that the components of the cavity fields stay finite as the system grows, then, the cavity field must be normalised by a quantity that is linear in the number of spin. Then we must have the following form of the cavity field
@@
Most importantly, we can explicitly compute the coefficients
$\hat{\gamma}_{\ell}=p_{\ell},\ \ \ \bar{\gamma}_{\ell}=2q_{\ell-1},\ \ \ \hat{\gamma}_{\ell}/\bar{\gamma}_{\ell}=p_{\ell}/2q_{\ell-1}\rightarrow 0$ (100)
@@
[…]
Notice in the last step we expanded the magnetisation states into the local magnetisations: we used a different superscript for the auxiliary variable $\ell^{*}$ to avoid confusion with the sub-index $\ell^{{}^{\prime}}$ before, that has quite different meaning. The free energy functional of the blanket is then
[…]
@@ 3.4 Example 2: Sherrington–Kirckpatrick model
this immediately imply the following form of the coefficients
$\hat{\gamma}_{\ell}=\sqrt{p_{\ell}},\ \ \ \bar{\gamma}_{\ell}=\sqrt{2q_{\ell-1}},\ \ \ \hat{\gamma}_{\ell}/\bar{\gamma}_{\ell}=\sqrt{p_{\ell}/2q_{\ell-1}}\rightarrow 0$ (107)
@@
$\hat{H}_{\ell}\left(\sigma_{V_{\ell}}\right)=\sqrt{p_{\ell}}\ H_{\mathrm{SK}}\left(\sigma_{V_{\ell}}\right),\ \ \ \bar{H}_{\ell}\left(\sigma_{V_{\ell}}|\sigma_{Q_{\ell-1}}\right)=\sigma_{V_{\ell}}\cdot\sqrt{2q_{\ell-1}}\ h_{V_{\ell}}\left(\sigma_{Q_{\ell-1}}\right)$ (108)
@@
$\sqrt{2q_{\ell-1}}\ h_{i}(\sigma_{Q_{\ell-1}})=\sum_{\ell^{*}<\ell}\sqrt{2p_{\ell^{*}}}\,\,h_{i}(\sigma_{V_{\ell^{*}}})$ (109)
@@
$\bar{f}_{\ell}\left(\sigma_{Q_{\ell-1}}\right)=-\frac{1}{\beta|V_{\ell}|}\sum_{i\in V_{\ell}}\log 2\cosh\,\beta\sqrt{2q_{\ell-1}}\,\,h_{i}(\sigma_{Q_{\ell-1}})=\\ =-\frac{1}{\beta|V_{\ell}|}\sum_{i\in V_{\ell}}\log 2\cosh\,\beta\sum_{\ell^{*}<\ell}\sqrt{2p_{\ell^{*}}}\,\,h_{i}(\sigma_{V_{\ell^{*}}})$ (110)
As said before, an alternative way to find the functional is the Cavity Method, [4, 33]: we showed a detailed derivation for the SK model in [1, 2, 3] (see also in Ref. [34] for the shortest derivation). Also, look at the Section 4 of [10], in Paragraph 4.2.1, for how to obtain the RSB theory as a special (stationary) case of the kernel picture, within the larger context of Lattice Field Theories [3, 10, 11, 19, 20].
@@ Acknowledgements
We are grateful to Pan Liming (USTC), Giampiero Bardella (Sapienza Universit Roma), and Nicola Kistler (Goethe University Frankfurt) for interesting discussions and suggestions. This research was partially funded by European Research Council (ERC), under the European Union’s Horizon 2020 research and innovation programme (Grant Agreement No. 694925).